# Stochastic Processes and Transport Plans

### v5

@@ 1 Introduction
Generative models with neural network approximations have shown impressive results in many applications, in particular in inverse problems and data assimilation. First developments in this direction were generative adversarial networks by Goodfellow et al. [17] in 2014 and variational autoencoders by Kingma and Welling [27] in 2013. In this paper, we are interested in recent flow matching techniques pioneered by Liu et al. [31, 32] in 2022/23 from the point of view of stochastic processes, by Lipman et al. [30] in 2023 via conditional probability paths and by Albergo et al. [2, 1] in 2022/2023 under the name of stochastic interpolants.. Flow matching is closely related to continuous normalizing flows (CNFs), also called neural ordinary differential equations by Chen et al. [11] in 2018, and to score-based diffusion models introduced by Sohl-Dickstein et al. [40] and Ermon and Song [42] in 2015 and 2019, respectively. The later ones can be considered as an instance of CNFs, but add the perspective of stochastic differential equations. To provide a more complete picture, we will briefly consider these approaches in Sections 10 and 11. Flow matching techniques stand out for their simple applicability and good scaling properties. They can be easily incorporated into other optimization techniques like plug-and-play approaches [33] and can be generalized for solving Bayesian inverse problems as we will see in Section 9. Recently, flow matching was generalized towards generator matching in [22, 24].
@@
Figure 1: Illustration of a curve from the standard Gaussian distribution to a Gaussian mixture in one dimension. The plot shows $t\mapsto\phi(t,x^{i})$, $i=1,2$ for two different samples $x^{i}$ (black, green), the vectors $(1,\partial_{t}\phi)$ and the red-blue color-coded velocity field $\partial_{t}\phi$. Courtesy: Blogpost [7]
Before learning the velocity fields via a neural network approximation, we have to determine appropriate curve-velocity pairs $(\mu_{t},v_{t})$ fulfilling in particular the continuity equation. In Sections 4 - 6, we provide three approaches. We will see that the first one follows as a special case both from the second and third one. Ultimately, we develop the second approach with Markov kernels to explain the setting of Lipman et al. in [30] also for measures without densities.
@@
    Curves via stochastic processes: Starting with a time differentiable stochastic process $(X_{t})_{t\in[0,1]}$, Liu et al. [31, 32] determined $(\mu_{t},v_{t})$ using the conditional expectation
@@
In the original paper [30], the notation ”conditional flow matching” (CFM) was used instead of FM which refers to a ”conditional flow path”. We prefer FM, since we will later deal with Bayesian inverse problems, where the notation of being ”conditional” is occupied in a different way. For the numerical simulation, we have on the one hand to sample from $\alpha_{t}=\mathcal{K}_{t}(y,\cdot)\times_{y}\mu_{1}$ and on the other hand to compute $v_{t}^{y}(x)$. Samples $y_{i}$, $i=1,\ldots,n$ from the data distribution $\mu_{1}$ are usually available and it should be easy to sample from the distributions $\mathcal{K}_{t}(y_{i},\cdot)$, $i=1,\ldots,n$. In the special case, where $(\nu_{t},v_{t})=(\mu_{t},v_{t})$ is induced by a coupling $\alpha$, we have a simple formula for $v_{t}^{y}$ for which the loss function becomes
@@ 2.1 Special Metric Spaces
Unfortunately, a sequence of probability measures $(\mu_{n})_{n}$ may not converge in the weak-* sense to a probability measure. For example, $\mu_{n}\coloneqq\delta_{n}$, $n\in\mathbb{N}$, converges weak-* towards $\mu\equiv 0$ which is not a probability measure. Therefore, the concept of narrow convergence222 In measure theory, narrow convergence of measures is also called weak convergence. We do not use the later notation to avoid confusion with the notation of weak convergence in functional analysis. was introduced: a sequence $(\mu_{n})_{n}\subset\mathcal{M}(\mathbb{R}^{d})$ converges narrowly to a measure $\mu\in\mathcal{M}(\mathbb{R}^{d})$, written $\mu_{n}\rightharpoonup\mu$ as $n\to\infty$, if
@@
If $\mu_{n}\in\mathcal{P}(\mathbb{R}^{d})$ with $\mu_{n}\rightharpoonup\mu$ as $n\to\infty$, then also $\mu\in\mathcal{P}(\mathbb{R}^{d})$. However, note that $\mathcal{M}(\mathbb{R}^{d})$ is not the dual space of $C_{b}(\mathbb{R}^{d})$, which is indeed the space $\text{rba}(\mathbb{R}^{d})$ of finitely additive set functions. For more information, see [36, Section 4.4].
@@
[…]
@@ 2.4 Couplings and Wasserstein Distance
The set $\Gamma(\mu,\nu)$ is a compact subset of $\mathcal{P}_{2}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ with respect to the narrow topology. The Wasserstein(-222) distance is defined by
@@
Indeed, the above minimum is attained, see, e.g. [38, Theorem 1.7], and the set of optimal plans is denoted by $\Gamma_{o}(\mu,\nu)$. The space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is a complete metric space. Speaking about $\mathcal{P}_{2}(\mathbb{R}^{d})$, we will always equip it with the $W_{2}$ metric. Then we have for a sequence $(\mu_{n})_{n}$ of measures in $\mathcal{P}_{2}(\mathbb{R}^{d})$ that
@@
see e.g. [35, Remark 2.31]. $\diamond$
@@ 3 Absolutely Continuous Curves in $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$
[4, Theorem 8.3.1] Let $\mu_{t}:I\to\mathcal{P}_{2}(\mathbb{R}^{d})$ be a narrowly continuous curve. Then $\mu_{t}$ is absolutely continuous if and only if there exists a Borel measurable vector field $v:I\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that the following two conditions are fulfilled:
@@
=$=\overline{\left\{\lambda(T-\operatorname{Id}):(\operatorname{Id},T)_{\#}\mu\in\Gamma_{o}(\mu,T_{\#}\mu),\;\lambda>0\right\}}^{L^{2}(\mathbb{R}^{d},\mu)},$ (26)
see [4, § 8]. The second description can be interpreted as locally moving mass in an ”optimal way”. Note that $\mathcal{T}_{\mu}$ is an infinite-dimensional subspace of $L^{2}(\mathbb{R}^{d},\mu)$ if $\mu$ is absolutely continuous, and it is just $\mathbb{R}^{d}$ if $\mu=\delta_{x}$, $x\in\mathbb{R}^{d}$. There is another description of $\mathcal{T}_{\mu}$, namely, for a vector field $v\in L^{2}(\mathbb{R}^{d},\mu)$, we have
@@
where the last equation is meant again in the distributional sense $\int_{\mathbb{R}^{d}}\langle w,\nabla\varphi\rangle\,\mathrm{d}\mu=0$ for all $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$, see [4, Lemma 8.4.2].
Now let $v_{t}\in\mathcal{T}_{\mu_{t}}$ be a vector field associated to $\mu_{t}$. For any other vector field $\tilde{v}_{t}$ associated to $\mu_{t}$, it holds $\nabla\cdot((v_{t}-\tilde{v}_{t})\mu)=0$, and hence by (27), $\|\tilde{v}_{t}\|_{L^{2}(\mathbb{R}^{d},\mu)}\geq\|v_{t}\|_{L^{2}(\mathbb{R}^{d},\mu)}$, meaning that $v_{t}$ has minimal $L^{2}(\mathbb{R}^{d},\mu_{t})$-norm. On the other hand, let $v_{t}$ be associated to $\mu_{t}$ having minimal norm. By [4, Theorem 8.3.1], there exists a vector field $\tilde{v}_{t}\in\mathcal{T}_{\mu_{t}}$ being associated to $\mu_{t}$ and having minimal norm. By the uniqueness shown above, it follows $v_{t}=\tilde{v}_{t}\in\mathcal{T}_{\mu_{t}}$.
@@
[4, Theorem 8.1.8] Let $\mu_{t}:I\to\mathcal{P}_{2}(\mathbb{R}^{d})$ be an absolutely continuous curve with associated vector field $v_{t}$ such that for every compact Borel set $B\subset\mathbb{R}^{d}$ it holds
@@
0 0 $=\varphi(b,\phi(b,x))-\varphi(a,\phi(a,x))=\int_{\mathbb{R}^{d}}\varphi(b,\phi(b,x))-\varphi(a,\phi(a,x))\mathrm{d}\mu_{0}$ (31)
@@ 4 Curves Induced by Couplings
In general, also optimal plans $\alpha\in\Gamma_{o}(\mu_{0},\mu_{1})$ may not be induced by a Monge map. However, if $\alpha$ is optimal and $\mu_{t}=e_{t,\sharp}\alpha$, then, for $s\in(0,1)$, there exists an optimal map $T^{s}$ such that $\Gamma_{o}(\mu_{s},\mu_{1})=\{(\operatorname{Id},T^{s})_{\sharp}\mu_{t}\}$, see [4, Lemma 7.2.1]. Moreover, by [4, Theorem 7.2.2], the curve induced by $T^{s}$ coincides, up to time parametrizations, with $\mu_{t}$ for $t\in[s,1]$, $s>0$. Figure 4(a) shows a curve from an optimal plan which does not arise from a map and Figure 4(b) a curve induced by a plan with Monge map. However, in each case, the plan $\Gamma_{o}(\mu_{s},\mu_{1})$ is induced by a map for $s\in(0,1)$.
@@
Moreover, curves induced by optimal plans are geodesics, see [4, Chapter 7.2].
@@
Then, by [3, Section 17], there exists a unique vector field $v\in L^{2}(e_{\sharp}\underline{{\bm{\alpha}}},\mathbb{R}^{d})$ such that
@@
Since every $e_{\sharp}{\bm{\alpha}}$-measurable function has a Borel measurable representative, see [8, Proposition 2.1.11], we can choose $v$ Borel measurable. We call a measurable vector field $v:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ induced by a plan $\alpha$, if it fulfills (40).
@@
Since the number of test functions $g_{n}$ is countable, there exists a zero set $\mathcal{I}\subset[0,1]$ such that (48) holds true on $[0,1]\setminus\mathcal{I}$ for all $n\in\mathbb{N}$. Since $\{g_{n}\}_{n\in\mathbb{N}}$ is dense in $C_{b}(\mathbb{R}^{d})$ we obtain that $v_{t}\mu_{t}=e_{t,\sharp}[(y-x)\alpha]$ for a.e. $t\in[0,1]$. More precisely, the last equation is an equation of vector-valued finite signed measures and thus it is enough to test equality at a dense subset of $C_{b}(\mathbb{R}^{d})$. ∎
@@
Assertion i) follows as in [3, Theorem 17.2, Lemma 17.3], see also [9, proof of Proposition 6].
@@
[…]
If $\alpha$ is not only optimal, but also induced by a Monge map $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$, it is often the case that the trajectories are straight, in the sense that the solution of $\partial_{t}\phi_{t}(x)=v_{t}(\phi_{t}(x)),\,\phi_{0}(x)=x$, is $T_{t}$. This is in particular the case, if $\mu_{0},\mu_{1}$ are empirical measures with the same number of points. For a proof and further cases, e.g. if both measures admit densities with a certain regularity, see [9, Proposition 16].
The above minimality property, i.e. laying in the tangent space, is quite special for optimal plans. Indeed, by the following example, this is not the case for arbitrary induced velocity fields, see also [31, Example 3.5].
@@
Hence, by (27), we know that every $u_{t}\in\mathcal{T}_{\mu_{t}}$ must satisfy $\int_{\mathbb{R}^{d}}\langle w,u_{t}\rangle\,\mathrm{d}\mu_{t}=0$ for a.e. $t\in[0,1]$. Unfortunately, the induced velocity field fulfills $\int_{\mathbb{R}^{d}}\langle w,v_{t}\rangle\,\mathrm{d}\mu_{t}>0$ for an non zero subset of [0,1][0,1][ 0 , 1 ] by the following reasons:
@@
Figure 6: Curve and vector field associated to $\alpha=\mu_{0}\times\mu_{1}$, where $\mu_{0}=\tfrac{1}{2}\delta_{x_{0}}+\tfrac{1}{2}\delta_{x_{1}}$ and $\mu_{1}=\tfrac{1}{3}\delta_{y_{0}}+\tfrac{2}{3}\delta_{y_{1}}$. Vectors are scaled by 0.20.20.2 for better visibility.
We have already seen that vector fields $v_{t}$ associated to optimal plans are minimal ones, meaning that $v_{t}\in\mathcal{T}_{\mu_{t}}$. This is in general not true for an independent coupling $\alpha=\mu_{0}\times\mu_{1}$ with an arbitrary $\mu_{0}$, see Figure 6. Fortunately, the independent coupling with a Gaussian marginal $\mu_{0}$ has this minimality property as the following example shows, see also [31, Section 5.1].

### v6

@@ 1 Introduction
Generative models with neural network approximations have shown impressive results in many applications, in particular in inverse problems and data assimilation. First developments in this direction were generative adversarial networks by Goodfellow et al. [goodfellow2014] in 2014 and variational autoencoders by Kingma and Welling [KW2013] in 2013. In this paper, we are interested in recent flow matching techniques pioneered by Liu et al. [liu2022rectified, liu2023flow] in 2022/23 from the point of view of stochastic processes, by Lipman et al. [lipman2023flow] in 2023 via conditional probability paths and by Albergo et al. [albergobuilding, albergo2023stochastic] in 2022/2023 under the name of stochastic interpolants.. Flow matching is closely related to continuous normalizing flows (CNFs), also called neural ordinary differential equations by Chen et al. [CRBD2018] in 2018, and to score-based diffusion models introduced by Sohl-Dickstein et al. [pmlr-v37-sohl-dickstein15] and Ermon and Song [SE2019] in 2015 and 2019, respectively. The later ones can be considered as an instance of CNFs, but add the perspective of stochastic differential equations. To provide a more complete picture, we will briefly consider these approaches in Sections 10 and 11. Flow matching techniques stand out for their simple applicability and good scaling properties. They can be easily incorporated into other optimization techniques like plug-and-play approaches [martin2024pnp] and can be generalized for solving Bayesian inverse problems as we will see in Section 9. Recently, flow matching was generalized towards generator matching in [holderrieth2025generator, JCHWS2025].
@@
Figure 1: Illustration of a curve from the standard Gaussian distribution to a Gaussian mixture in one dimension. The plot shows $t\mapsto\phi(t,x^{i})$, $i=1,2$ for two different samples $x^{i}$ (black, green), the vectors $(1,\partial_{t}\phi)$ and the red-blue color-coded velocity field $\partial_{t}\phi$. Courtesy: Blogpost [sego]
Before learning the velocity fields via a neural network approximation, we have to determine appropriate curve-velocity pairs $(\mu_{t},v_{t})$ fulfilling in particular the continuity equation. In Sections 4 - 6, we provide three approaches. We will see that the first one follows as a special case both from the second and third one. Ultimately, we develop the second approach with Markov kernels to explain the setting of Lipman et al. in [lipman2023flow] also for measures without densities.
@@
    Curves via stochastic processes: Starting with a time differentiable stochastic process $(X_{t})_{t\in[0,1]}$, Liu et al. [liu2022rectified, liu2023flow] determined $(\mu_{t},v_{t})$ using the conditional expectation
@@
In the original paper [lipman2023flow], the notation ”conditional flow matching” (CFM) was used instead of FM which refers to a ”conditional flow path”. We prefer FM, since we will later deal with Bayesian inverse problems, where the notation of being ”conditional” is occupied in a different way. For the numerical simulation, we have on the one hand to sample from $\alpha_{t}=\mathcal{K}_{t}(y,\cdot)\times_{y}\mu_{1}$ and on the other hand to compute $v_{t}^{y}(x)$. Samples $y_{i}$, $i=1,\ldots,n$ from the data distribution $\mu_{1}$ are usually available and it should be easy to sample from the distributions $\mathcal{K}_{t}(y_{i},\cdot)$, $i=1,\ldots,n$. In the special case, where $(\nu_{t},v_{t})=(\mu_{t},v_{t})$ is induced by a coupling $\alpha$, we have a simple formula for $v_{t}^{y}$ for which the loss function becomes
@@ 2.1 Special Metric Spaces
Unfortunately, a sequence of probability measures $(\mu_{n})_{n}$ may not converge in the weak-* sense to a probability measure. For example, $\mu_{n}\coloneqq\delta_{n}$, $n\in\mathbb{N}$, converges weak-* towards $\mu\equiv 0$ which is not a probability measure. Therefore, the concept of narrow convergence [^2] In measure theory, narrow convergence of measures is also called weak convergence. We do not use the later notation to avoid confusion with the notation of weak convergence in functional analysis. was introduced: a sequence $(\mu_{n})_{n}\subset\mathcal{M}(\mathbb{R}^{d})$ converges narrowly to a measure $\mu\in\mathcal{M}(\mathbb{R}^{d})$, written $\mu_{n}\rightharpoonup\mu$ as $n\to\infty$, if
@@
If $\mu_{n}\in\mathcal{P}(\mathbb{R}^{d})$ with $\mu_{n}\rightharpoonup\mu$ as $n\to\infty$, then also $\mu\in\mathcal{P}(\mathbb{R}^{d})$. However, note that $\mathcal{M}(\mathbb{R}^{d})$ is not the dual space of $C_{b}(\mathbb{R}^{d})$, which is indeed the space $\text{rba}(\mathbb{R}^{d})$ of finitely additive set functions. For more information, see [PPST2023, Section 4.4].
@@
where $\|\cdot\|$ is the Euclidean norm on $\mathbb{R}^{m}$. In case of the Lebesgue measure $\mu=\mathcal{L}$, we will skip the $\mu$ in the notation. Furthermore, we write $L^{1}$ for $L^{1}(\mathbb{R},\mathcal{L})$.
@@ 2.4 Couplings and Wasserstein Distance
The set $\Gamma(\mu,\nu)$ is a compact subset of $\mathcal{P}_{2}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ with respect to the narrow topology. The Wasserstein(-$2$) distance is defined by
@@
Indeed, the above minimum is attained, see, e.g. [santambrogio2015optimal, Theorem 1.7], and the set of optimal plans is denoted by $\Gamma_{o}(\mu,\nu)$. The space $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ is a complete metric space. Speaking about $\mathcal{P}_{2}(\mathbb{R}^{d})$, we will always equip it with the $W_{2}$ metric. Then we have for a sequence $(\mu_{n})_{n}$ of measures in $\mathcal{P}_{2}(\mathbb{R}^{d})$ that
@@
see e.g. [peyre2019computational, Remark 2.31]. $\diamond$
@@ 3 Absolutely Continuous Curves in $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$
[AGS2008, Theorem 8.3.1] Let $\mu_{t}:I\to\mathcal{P}_{2}(\mathbb{R}^{d})$ be a narrowly continuous curve. Then $\mu_{t}$ is absolutely continuous if and only if there exists a Borel measurable vector field $v:I\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that the following two conditions are fulfilled:
@@
$=\overline{\left\{\lambda(T-\operatorname{Id}):(\operatorname{Id},T)_{\#}\mu\in\Gamma_{o}(\mu,T_{\#}\mu),\;\lambda>0\right\}}^{L^{2}(\mathbb{R}^{d},\mu)},$ (26)
see [AGS2008, § 8]. The second description can be interpreted as locally moving mass in an ”optimal way”. Note that $\mathcal{T}_{\mu}$ is an infinite-dimensional subspace of $L^{2}(\mathbb{R}^{d},\mu)$ if $\mu$ is absolutely continuous, and it is just $\mathbb{R}^{d}$ if $\mu=\delta_{x}$, $x\in\mathbb{R}^{d}$. There is another description of $\mathcal{T}_{\mu}$, namely, for a vector field $v\in L^{2}(\mathbb{R}^{d},\mu)$, we have
@@
where the last equation is meant again in the distributional sense $\int_{\mathbb{R}^{d}}\langle w,\nabla\varphi\rangle\,\mathrm{d}\mu=0$ for all $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$, see [AGS2008, Lemma 8.4.2].
Now let $v_{t}\in\mathcal{T}_{\mu_{t}}$ be a vector field associated to $\mu_{t}$. For any other vector field $\tilde{v}_{t}$ associated to $\mu_{t}$, it holds $\nabla\cdot((v_{t}-\tilde{v}_{t})\mu)=0$, and hence by (27), $\|\tilde{v}_{t}\|_{L^{2}(\mathbb{R}^{d},\mu)}\geq\|v_{t}\|_{L^{2}(\mathbb{R}^{d},\mu)}$, meaning that $v_{t}$ has minimal $L^{2}(\mathbb{R}^{d},\mu_{t})$-norm. On the other hand, let $v_{t}$ be associated to $\mu_{t}$ having minimal norm. By [AGS2008, Theorem 8.3.1], there exists a vector field $\tilde{v}_{t}\in\mathcal{T}_{\mu_{t}}$ being associated to $\mu_{t}$ and having minimal norm. By the uniqueness shown above, it follows $v_{t}=\tilde{v}_{t}\in\mathcal{T}_{\mu_{t}}$.
@@
[AGS2008, Theorem 8.1.8] Let $\mu_{t}:I\to\mathcal{P}_{2}(\mathbb{R}^{d})$ be an absolutely continuous curve with associated vector field $v_{t}$ such that for every compact Borel set $B\subset\mathbb{R}^{d}$ it holds
@@
$0$ $=\varphi(b,\phi(b,x))-\varphi(a,\phi(a,x))=\int_{\mathbb{R}^{d}}\varphi(b,\phi(b,x))-\varphi(a,\phi(a,x))\mathrm{d}\mu_{0}$ (31)
@@ 4 Curves Induced by Couplings
In general, also optimal plans $\alpha\in\Gamma_{o}(\mu_{0},\mu_{1})$ may not be induced by a Monge map. However, if $\alpha$ is optimal and $\mu_{t}=e_{t,\sharp}\alpha$, then, for $s\in(0,1)$, there exists an optimal map $T^{s}$ such that $\Gamma_{o}(\mu_{s},\mu_{1})=\{(\operatorname{Id},T^{s})_{\sharp}\mu_{t}\}$, see [AGS2008, Lemma 7.2.1]. Moreover, by [AGS2008, Theorem 7.2.2], the curve induced by $T^{s}$ coincides, up to time parametrizations, with $\mu_{t}$ for $t\in[s,1]$, $s>0$. Figure 4(a) shows a curve from an optimal plan which does not arise from a map and Figure 4(b) a curve induced by a plan with Monge map. However, in each case, the plan $\Gamma_{o}(\mu_{s},\mu_{1})$ is induced by a map for $s\in(0,1)$.
@@
Moreover, curves induced by optimal plans are geodesics, see [AGS2008, Chapter 7.2].
@@
Then, by [ambrosio2021lectures, Section 17], there exists a unique vector field $v\in L^{2}(e_{\sharp}\underline{{\bm{\alpha}}},\mathbb{R}^{d})$ such that
@@
Since every $e_{\sharp}{\bm{\alpha}}$-measurable function has a Borel measurable representative, see [bogachev2007measure, Proposition 2.1.11], we can choose $v$ Borel measurable. We call a measurable vector field $v:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ induced by a plan $\alpha$, if it fulfills (40).
@@
Since the number of test functions $g_{n}$ is countable, there exists a zero set $\mathcal{I}\subset[0,1]$ such that (48) holds true on $0,1]\setminus\mathcal{I}$ for all $n\in\mathbb{N}$. Since $\{g_{n}\}_{n\in\mathbb{N}}$ is dense in $C_{b}(\mathbb{R}^{d})$ we obtain that $v_{t}\mu_{t}=e_{t,\sharp}[(y-x)\alpha]$ for a.e. $t\in[0,1]$. More precisely, the last equation is an equation of vector-valued finite signed measures and thus it is enough to test equality at a dense subset of $C_{b}(\mathbb{R}^{d})$. ∎
@@
Assertion i) follows as in [ambrosio2021lectures, Theorem 17.2, Lemma 17.3], see also [chemseddine2024conditional, proof of Proposition 6].
@@
By Theorem 4.6ii)ii) we know that $\|v_{t}\|_{L}^{2}(\mathbb{R}^{d},\mu_{t})=|\mu^{\prime}_{t}|$ for a.e. $t\in[0,1]$. Thus [AGS2008, Proposition 8.4.5] implies $v_{t}\in\mathcal{T}_{\mu_{t}}$ for a.e. $t\in[0,1]$. ∎
If $\alpha$ is not only optimal, but also induced by a Monge map $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$, it is often the case that the trajectories are straight, in the sense that the solution of $\partial_{t}\phi_{t}(x)=v_{t}(\phi_{t}(x)),\,\phi_{0}(x)=x$, is $T_{t}$. This is in particular the case, if $\mu_{0},\mu_{1}$ are empirical measures with the same number of points. For a proof and further cases, e.g. if both measures admit densities with a certain regularity, see [chemseddine2024conditional, Proposition 16].
The above minimality property, i.e. laying in the tangent space, is quite special for optimal plans. Indeed, by the following example, this is not the case for arbitrary induced velocity fields, see also [liu2022rectified, Example 3.5].
@@
Hence, by (27), we know that every $u_{t}\in\mathcal{T}_{\mu_{t}}$ must satisfy $\int_{\mathbb{R}^{d}}\langle w,u_{t}\rangle\,\mathrm{d}\mu_{t}=0$ for a.e. $t\in[0,1]$. Unfortunately, the induced velocity field fulfills $\int_{\mathbb{R}^{d}}\langle w,v_{t}\rangle\,\mathrm{d}\mu_{t}>0$ for an non zero subset of $[0,1]$ by the following reasons:
@@
Figure 6: Curve and vector field associated to $\alpha=\mu_{0}\times\mu_{1}$, where $\mu_{0}=\tfrac{1}{2}\delta_{x_{0}}+\tfrac{1}{2}\delta_{x_{1}}$ and $\mu_{1}=\tfrac{1}{3}\delta_{y_{0}}+\tfrac{2}{3}\delta_{y_{1}}$. Vectors are scaled by $0.2$ for better visibility.
We have already seen that vector fields $v_{t}$ associated to optimal plans are minimal ones, meaning that $v_{t}\in\mathcal{T}_{\mu_{t}}$. This is in general not true for an independent coupling $\alpha=\mu_{0}\times\mu_{1}$ with an arbitrary $\mu_{0}$, see Figure 6. Fortunately, the independent coupling with a Gaussian marginal $\mu_{0}$ has this minimality property as the following example shows, see also [liu2022rectified, Section 5.1].